OpenStudy (anonymous):

FAN & MEDAL! :) PLEASE HELP ME! The perimeter of a rectangle is 120 feet. The ratio of the width to the length is 2:3. Find the Length and the width.

OpenStudy (anonymous):

Length = 12 Width = 8

OpenStudy (anonymous):

We know that the length is 3 times x and the width is 2 times x. The perimeter is the two long sides added to the two widths. So we have 3x+3x+2x+2x=40. Collect terms gives 10x = 40. x = 4 So the length is = 3*4 = 12 and the width = 2*4 = 8